# Understanding Flow Behaviors of Supercooled Liquids by Embodying Solid-Liquid Duality at Particle Level

### v1

@@ II.4 Functional form of $\tau_{\mathrm{sh}}(\tau_{\mathrm{LC}})$
Figure 11: (a) and (b) show the fitted $\tau_{0}$ as a function of $\dot{\gamma}^{-1}$ for sample A and sample B, respectively. Dashed lines in (a) and (b) denote the linear relation between $\tau_{0}$ and $\dot{\gamma}^{-1}$. (c) and (d) show the fitted $\tau_{\text{sh}}^{\infty}$ as a function of $\dot{\gamma}$ for sample A and sample B, respectively. Dashed lines in (c) and (d) denote the relation of $\tau_{\mathrm{sh}}^{\infty}\sim\dot{\gamma}^{-\lambda_{\tau}}$. (e) and (f) give the rescaled relations of $\tau_{\mathrm{sh}}(\tau_{\mathrm{LC}})$, denoted as $\tau_{\mathrm{sh}}^{*}(\tau_{\mathrm{LC}}^{*})$, for sample A and sample B, respectively. Dashed lines in (e) and (f) represent the master curves given by Eq. 4.
@@
where $\tau_{\mathrm{sh}}^{\infty}$ and $\tau_{0}$ are fitting parameters. $\tau_{0}$ governs the position at which $\tau_{\mathrm{sh}}(\tau_{\mathrm{LC}})$ significantly deviates from $\tau_{\mathrm{sh}}=\tau_{\mathrm{LC}}$. It reflects the time scale of the external shear. Thus, we expect that $\tau_{0}$ depends on $\dot{\gamma}^{-1}$ linearly. Figure 11 (a) and (b) give the fitted $\tau_{0}$ as a function of $\dot{\gamma}^{-1}$ for sample A and sample B, respectively, from where the linearity is clearly seen. $\tau_{\mathrm{sh}}^{\infty}$ represents the dynamics of the slowest particles in the flow. These particles, according to our above analysis, are mostly involved in LERs. We can define the characteristic strain for these particles by $\gamma_{\mathrm{sh}}^{\infty}=\dot{\gamma}\tau_{\mathrm{sh}}^{\infty}$. In Ref. [42], it is demonstrated that in the shear-thinning regime, the strain of LER $\gamma_{\mathrm{LER}}$ varies with $\dot{\gamma}$ by $\gamma_{\mathrm{LER}}\sim\dot{\gamma}^{\epsilon}$ with $\epsilon=1-\lambda$. Therefore, we expect $\gamma_{\mathrm{sh}}^{\infty}\sim\dot{\gamma}^{\epsilon}$, or equivalently, $\tau_{\mathrm{sh}}^{\infty}\sim\dot{\gamma}^{-\lambda}$. The fitted $\tau_{\mathrm{sh}}^{\infty}$ as a function of $\dot{\gamma}$ are plotted in log-log scale for both samples in Fig. 11 (c) and (d). The relation of $\tau_{\mathrm{sh}}^{\infty}(\dot{\gamma})$ indeed exhibits a power-law dependence, denoted as $\tau_{\mathrm{sh}}^{\infty}\sim\dot{\gamma}^{-\lambda_{\tau}}$. The fitted value of the power $\lambda_{\tau}$ is $0.70$ for sample A, and $0.65$ for sample B. As we expect, these values are very close to the corresponding $\lambda$. Note that the power law of $\tau_{\mathrm{sh}}^{\infty}\sim\dot{\gamma}^{-\lambda_{\tau}}$ covers the whole range of $\dot{\gamma}$ studied in this work, and is not limited to the shear-thinning regime. According to these considerations, we can rescale $\tau_{\mathrm{sh}}$ as $\tau_{\mathrm{sh}}^{*}=\tau_{\mathrm{sh}}/(c\,\dot{\gamma}^{-\lambda})$, and rescale $\tau_{\mathrm{LC}}$ as $\tau_{\mathrm{LC}}^{*}=\tau_{\mathrm{LC}}/(a\dot{\gamma}^{-1}+b)$, where $a$, $b$ and $c$ are material-dependent parameters. Then, Eq. 3 is rewritten as
@@
We rescale our data found from different conditions with the above method for both samples, and show the results in Fig. 11 (e) and (f). All data well collapse onto the master curve given by Eq. 4. Note that, since the exponent $\lambda$ is involved in this rescaling, Eq. 4 implies an intrinsic relation between the nonlinear rheology of the system and the interplay between convection and thermal activation.
Here, both the values of $\lambda$ of two samples are close to $2/3$. Similar results are also observed in other LJ supercooled liquids in a wide range of supercooled temperatures [89]. Moreover, for Brownian dynamics simulations of supercooled charged colloids with hard-sphere Yukawa interaction, it is found that $\lambda$ ranges from $0.64$ to $0.7$, which are also closely around $2/3$ [42]. In fact, $2/3$ is the prediction of a mean-field theory [90]. Therefore, by assuming $\tau_{\mathrm{sh}}^{\infty}=c\,\dot{\gamma}^{-2/3}$, the obtained $\tau_{\mathrm{sh}}^{\infty}$ still effectively rescale the data of $\tau_{\mathrm{sh}}(\tau_{\mathrm{LC}})$ onto Eq. 4.
We will get back to the form of $\tau_{\mathrm{sh}}(\tau_{\mathrm{LC}})$ in Sec. V.
@@ III SOLID-LIQUID DUALITY AND RHEOLOGY
In the preceding section, we establish the correlation between DH and local mechanical response to imposed shear in supercooled liquids. It is found that the way of local response highly depends on the comparison between the local mobility, represented by $\tau_{\mathrm{LC}}$, and the time scale of external shear, which linearly depends on $\dot{\gamma}^{-1}$. Though the spectrum of $\tau_{\mathrm{LC}}$ is smooth, the local response exhibits two major ways: solid-like for slow particles, and liquid-like for fast particles. This solid-liquid duality in response can be viewed as a specific implementation of Maxwell’s insight into viscoelasticity at the microscopic level. Additionally, it is crucial to note that slow/fast particles do not randomly distribute through the system. Instead, they tend to aggregate into clusters. With this picture, we will explore the microscopic mechanisms of two important rheological behaviors for supercooled liquids, namely, steady shear and start-up shear, in this section.
@@ III.1 Rearrangements of solid-like and liquid-like regions
To employ the microscopic solid-liquid duality, it is important to clarify the difference between the rearrangements of solid-like and liquid-like regions first. Figure 12 (a) and (b) display the stress evolutions of a solid-like region (the LER L5 in Fig. 10) and a liquid-like region (the fast cluster F1 in Fig. 10) in simple A, respectively. For each region, we calculate the ratio of constituent particles that have undergone cage jump since $t=t_{\mathrm{in}}$, denoted as $R_{\mathrm{cj}}(t)$. Here, $t_{\mathrm{in}}$, marked in Fig. 12 (a) and (b), is the beginning time of the elastic deformation of LER L5. The results of $R_{\mathrm{cj}}(t)$ are shown in Fig. 12 (c) and (d). As seen from Fig. 12 (c), for the solid-like region, $R_{\mathrm{cj}}(t)$ increases slowly in the elastic deformation stage, and then grows much more rapidly as the cluster enters the plastic yielding process. In contrast, $R_{\mathrm{cj}}(t)$ for the liquid-like region increases in a nearly linear form, as shown in Fig. 12 (d). The difference between Fig. 12 (c) and (d) reflect that solid-like region and liquid-like region rearrange in fundamentally different ways, as we will discuss in following two paragraphs.
Figure 12: (a) and (b) reproduce the stress evolutions of the LER L5 (representing a solid-like region) and the fast cluster F1 (representing a liquid-like region) in Fig. 10. $t_{\mathrm{in}}$ is the time at which the LER L5 starts to accumulate elastic stress. (c) and (d) give the jump ratio $R_{\mathrm{cj}}(t)$ for the above two regions. In (a) and (c), we denote the end of elastic deformation stage by a vertical dashed line.
@@
For solid-like regions, $R_{\mathrm{cj}}(t)$ exhibits an evident crossover at the end of the elastic deformation stage. To understand this nontrivial phenomenon, we recall that the yielding process of amorphous solids is collective and hierarchical [59], where the facilitation due to the long-ranged anisotropic elasticity-mediated interaction plays a crucial role [92, 93, 94]. Particularly, when only a small fraction of particles undergo rearrangements, the system remarkably loses its ability to further accumulate elastic stress: As suggested in Ref. [95], the failure of amorphous solids occurs when only about $10\%$ of particles are involved in STZs. As for LER, its response is very similar to that of amorphous solids, as shown in Fig. 10 and Ref. [83]. Especially, the end of its elastic deformation is accompanied by the emergence of cage-jump spots that manifest STZs [83]. With these considerations, the observation shown in Fig. 12 (c) is understandable: During the elastic deformation stage of LER, cage jumps are relatively rare, and $R_{\mathrm{cj}}(t)$ grows slowly with $t$. As $R_{\mathrm{cj}}(t)$ exceeds about $10\%$, the elasticity cannot be well sustained any longer. Then, plastic yielding and disintegration successively occur, accompanied by overwhelming cage jumps that result in the rapid growth of $R_{\mathrm{cj}}(t)$.
@@ III.2 Steady shear
Under steady shear, both samples exhibit shear thinning, as shown in Fig. 1. It can be described by a phenomenological expression $\eta(\dot{\gamma})\approx\eta_{\mathrm{eq}}\big{/}\left[1+(\tau_{\eta}\dot{\gamma})^{\lambda}\right]$, where $\eta_{\mathrm{eq}}$ is the zero-shear viscosity, and $\tau_{\eta}$ is a parameter with the dimension of time [2, 43, 96]. As $\dot{\gamma}$ increases from null, $\eta$ gradually deviates from $\eta_{\mathrm{eq}}$ and finally behaves with a power law $\eta(\dot{\gamma})\sim\dot{\gamma}^{-\lambda}$. To understand this complicated nonlinear behavior, we employ the solid-liquid duality by expressing the shear stress $\sigma$ as:
@@
where the first term in the right-hand side of Eq. 5 represents the contribution from slow particles, and the second term represents the contribution from fast particles. We discuss the former first. $p_{\mathrm{slow}}$ is the fraction of slow particles. As demonstrated in Sec. II.3, most slow particles are involved in the elastic stage of LERs. Thus, this contribution is given by the average elastic stress of LER. Here, $G_{\mathrm{eff}}$ is the effective shear modulus of LER, and $\tau_{\mathrm{el}}$ is the average time of duration of the elastic deformation of LER. In flow, an LER undergoes elastic loading, and then a sharp disintegration to unload. The factor $1/2$ accounts for the averaging effect of this saw-tooth-like loading-unloading process.
$\tau_{\mathrm{el}}$ and $G_{\mathrm{eff}}$ are needed to quantify the contribution from slow particles. $\tau_{\mathrm{el}}$ is evaluated as follows. As we discuss in the preceding subsection, the end of the elastic deformation stage of an LER is accompanied by the cage jumps of about $10\%$ of its constituent particles. Therefore, we determine $\tau_{\mathrm{el}}$ as the waiting time that $10\%$ of slow particles have relaxed according to the spectrum of $\tau_{\mathrm{sh}}$.
We now turn our attention to the effective modulus of LER $G_{\mathrm{eff}}$. By “effective”, we mean that it encapsulates both elastic and plastic effects [71, 97]. To be specific, even in the elastic stage of LER, small amount of nonaffine displacements of constituent particles take place, as shown in Fig. 12 (c) and Ref. [83]. This plasticity slightly weakens the modulus of LER. Therefore, the modulus during the elastic stage can be written as $G_{\mathrm{eff}}=wG_{\mathrm{ini}}$, where $G_{\mathrm{ini}}$ is the initial shear modulus of LER, and $w$ ($w<1$) is a weakening factor reflecting this plasticity. To estimate $w$, we adopt a strategy that ties the variation in shear modulus to the neighbor loss process [71, 97]. The detail for calculating $w$ and $G_{\mathrm{eff}}$ is given in Appendix B. For all studied conditions, we find that $G_{\mathrm{eff}}\approx 0.8G_{\mathrm{ini}}$.
For fast particles, as we show in Fig. 10, they do not exhibit elastic deformation. Instead, such regions behave like viscous liquids. Considering that the average persistence time is found to be proportional to viscosity [62], we can estimate the viscosity of fast particles $\eta_{\mathrm{fast}}$ by the relation:
@@
where $\bar{\tau}_{\mathrm{LC,eq}}$ is the average of $\tau_{\mathrm{LC}}$ at equilibrium, and $\bar{\tau}_{\mathrm{sh,fast}}$ is the average of $\tau_{\mathrm{sh}}$ of fast particles at the given shear rate. Here, $\bar{\tau}_{\mathrm{sh}}$ can be approximated by $\bar{\tau}_{\mathrm{LC}}$ for fast particles. With Eq. 6, we obtain the second term in the right-hand side of Eq. 5.
Figure 13: In both (a) (representing sample A) and (b) (representing sample B), the viscosity estimated by Eq. 5, the viscosity directly measured from MD data, and the proportion of slow particles $p_{\mathrm{slow}}$ are given.
Figure 13 shows the results of viscosity calculated by Eq. 5 for two samples. The accurate results, shown in Fig. 1, are replotted here for comparison. The agreement between two results is remarkable. We also show the results of $p_{\mathrm{slow}}$ in Fig. 13. It is seen that $p_{\mathrm{slow}}$ is negligible at the Newtonian regime. As $\dot{\gamma}$ increases, $p_{\mathrm{slow}}$ becomes finite, meanwhile, noticeable shear thinning appears. As $\dot{\gamma}$ further increases, $p_{\mathrm{slow}}$ grows rapidly, and then saturates in the shear thinning regime. These observations clearly reveal the relation between DH and the onset of shear thinning: Due to the heterogeneous dynamics, different regions in a supercooled liquid respond to shear in different ways. Slow particles tend to aggregate into LERs. The accumulation of elastic stress and strain of LER, which is limited by thermal activation and shear-driven effects, cannot catch up with the growth of $\dot{\gamma}$. Thus, the viscosity contributed by slow particles shear-thins. As $\dot{\gamma}$ becomes larger, more particles are involved as slow particles, results in progressively prominent shear thinning. When $\dot{\gamma}$ is very large, most particles are involved in LERs. Consequently, the nonlinear rheology of the system, described by $\eta\propto\dot{\gamma}^{-\lambda}$, is dominated by the behavior of LER. As we demonstrate in Ref. [42], in the shear-thinning regime, the characteristic length $\xi_{\mathrm{LER}}$ and strain $\gamma_{\mathrm{LER}}$ of LER, respectively, depend on $\dot{\gamma}$ by $\xi_{\mathrm{LER}}\propto\dot{\gamma}^{-v}$ and $\gamma_{\mathrm{LER}}\propto\dot{\gamma}^{\epsilon}$. Then, the exponent $\lambda$ is determined by LER through the scaling relation $\lambda=4v=1-\epsilon$.

### v2

@@ II.4 Functional form of $\tau_{\mathrm{sh}}(\tau_{\mathrm{LC}})$
Figure 11: (a) and (b) show the fitted $\tau_{0}$ as a function of $\dot{\gamma}^{-1}$ for sample A and sample B, respectively. Dashed lines in (a) and (b) denote the linear relation between $\tau_{0}$ and $\dot{\gamma}^{-1}$. (c) and (d) show the fitted $\tau_{\text{sh}}^{\infty}$ as a function of $\dot{\gamma}$ for sample A and sample B, respectively. Dashed lines in (c) and (d) denote the relation of $\tau_{\mathrm{sh}}^{\infty}\sim\dot{\gamma}^{-\lambda_{\tau}}$. (e) and (f) give the rescaled relations of $\tau_{\mathrm{sh}}(\tau_{\mathrm{LC}})$, denoted as $\tau_{\mathrm{sh}}^{*}(\tau_{\mathrm{LC}}^{*})$, for sample A and sample B, respectively. Dashed lines in (e) and (f) represent the master curves given by equation 4.
@@
where $\tau_{\mathrm{sh}}^{\infty}$ and $\tau_{0}$ are fitting parameters. $\tau_{0}$ governs the position at which $\tau_{\mathrm{sh}}(\tau_{\mathrm{LC}})$ significantly deviates from $\tau_{\mathrm{sh}}=\tau_{\mathrm{LC}}$. It reflects the time scale of the external shear. Thus, we expect that $\tau_{0}$ depends on $\dot{\gamma}^{-1}$ linearly. Figure 11 (a) and (b) give the fitted $\tau_{0}$ as a function of $\dot{\gamma}^{-1}$ for sample A and sample B, respectively, from where the linearity is clearly seen. $\tau_{\mathrm{sh}}^{\infty}$ represents the dynamics of the slowest particles in the flow. These particles, according to our above analysis, are mostly involved in LERs. We can define the characteristic strain for these particles by $\gamma_{\mathrm{sh}}^{\infty}=\dot{\gamma}\tau_{\mathrm{sh}}^{\infty}$. In Ref. [44], it is demonstrated that in the shear-thinning regime, the strain of LER $\gamma_{\mathrm{LER}}$ varies with $\dot{\gamma}$ by $\gamma_{\mathrm{LER}}\sim\dot{\gamma}^{\epsilon}$ with $\epsilon=1-\lambda$. Therefore, we expect $\gamma_{\mathrm{sh}}^{\infty}\sim\dot{\gamma}^{\epsilon}$, or equivalently, $\tau_{\mathrm{sh}}^{\infty}\sim\dot{\gamma}^{-\lambda}$. The fitted $\tau_{\mathrm{sh}}^{\infty}$ as a function of $\dot{\gamma}$ are plotted in log-log scale for both samples in figure 11 (c) and (d). The relation of $\tau_{\mathrm{sh}}^{\infty}(\dot{\gamma})$ indeed exhibits a power-law dependence, denoted as $\tau_{\mathrm{sh}}^{\infty}\sim\dot{\gamma}^{-\lambda_{\tau}}$. The fitted value of the power $\lambda_{\tau}$ is $0.70$ for sample A, and $0.65$ for sample B. As we expect, these values are very close to the corresponding $\lambda$. Note that the power law of $\tau_{\mathrm{sh}}^{\infty}\sim\dot{\gamma}^{-\lambda_{\tau}}$ covers the whole range of $\dot{\gamma}$ studied in this work, and is not limited to the shear-thinning regime. According to these considerations, we can rescale $\tau_{\mathrm{sh}}$ as $\tau_{\mathrm{sh}}^{*}=\tau_{\mathrm{sh}}/(c\,\dot{\gamma}^{-\lambda})$, and rescale $\tau_{\mathrm{LC}}$ as $\tau_{\mathrm{LC}}^{*}=\tau_{\mathrm{LC}}/(a\dot{\gamma}^{-1}+b)$, where $a$, $b$ and $c$ are material-dependent parameters. Then, equation 3 is rewritten as
@@
We rescale our data found from different conditions with the above method for both samples, and show the results in figure 11 (e) and (f). All data well collapse onto the master curve given by equation 4.
Both the values of $\lambda$ of two samples are close to $2/3$. Similar results are also observed in other LJ liquids in a wide range of supercooled temperatures [89]. In fact, $2/3$ is the prediction of a mean-field theory [90]. Recently, we furtherly show that the value of $\lambda$ is determined by the behaviors of localized elasticity and slightly alters among different samples [44, 83]. Nevertheless, by assuming $\tau_{\mathrm{sh}}^{\infty}=c\,\dot{\gamma}^{-2/3}$, the obtained $\tau_{\mathrm{sh}}^{\infty}$ still effectively collapse the data of $\tau_{\mathrm{sh}}(\tau_{\mathrm{LC}})$ with equation 4.
We will provide a model analysis on the form of $\tau_{\mathrm{sh}}(\tau_{\mathrm{LC}})$ in section V.
@@ III SOLID-LIQUID DUALITY AND RHEOLOGY
In the preceding section, we establish the correlation between DH and local mechanical response to imposed shear in supercooled liquids. It is found that the way of local response highly depends on the comparison between the local mobility, represented by $\tau_{\mathrm{LC}}$, and the time scale of external shear, which linearly depends on $\dot{\gamma}^{-1}$. Though the distribution of $\tau_{\mathrm{LC}}$ value is smooth, the local response exhibits two major ways: solid-like for slow particles, and liquid-like for fast particles. This solid-liquid duality in response can be viewed as a specific implementation of Maxwell’s insight into viscoelasticity at the microscopic level. Additionally, it is crucial to note that slow/fast particles do not randomly distribute through the system. Instead, they tend to aggregate into clusters. With this picture, we will explore the microscopic mechanisms of two important rheological behaviors for supercooled liquids, namely, steady shear and start-up shear, in this section.
@@ III.1 Rearrangements of solid-like and liquid-like regions
To employ the microscopic solid-liquid duality, it is important to clarify the difference between the rearrangements of solid-like and liquid-like regions first. Figure 12 (a) and (b) display the stress evolutions of a solid-like region (the LER L5 in figure 10) and a liquid-like region (the fast cluster F1 in figure 10) in simple A, respectively. For each region, we calculate the ratio of constituent particles that have undergone cage jump since $t=t_{\mathrm{in}}$, denoted as $R_{\mathrm{cj}}(t)$. Here, $t_{\mathrm{in}}$, marked in figure 12 (a) and (b), is the beginning time of the elastic deformation of LER L5. The results of $R_{\mathrm{cj}}(t)$ are shown in figure 12 (c) and (d). As seen from figure 12 (c), for the solid-like region, $R_{\mathrm{cj}}(t)$ increases slowly in the elastic deformation stage, and then grows much more rapidly as the cluster enters the plastic yielding process. In contrast, $R_{\mathrm{cj}}(t)$ for the liquid-like region increases in a nearly linear form, as shown in figure 12 (d). The difference between figure 12 (c) and (d) reflect that solid-like region and liquid-like region rearrange in fundamentally different ways, as we will discuss in following two paragraphs.
Figure 12: (a) and (b) reproduce the stress evolutions of the LER L5 (representing a solid-like region) and the fast cluster F1 (representing a liquid-like region) in figure 10. $t_{\mathrm{in}}$ is the time at which the LER L5 starts to accumulate elastic stress. (c) and (d) give the jump ratio $R_{\mathrm{cj}}(t)$ for the above two regions. In (a) and (c), we denote the end of elastic deformation stage by a vertical dashed line.
@@
For solid-like regions, $R_{\mathrm{cj}}(t)$ exhibits an evident crossover at the end of the elastic deformation stage. To understand this nontrivial phenomenon, we recall that the yielding process of amorphous solids is collective and hierarchical [59], where the facilitation due to the long-ranged anisotropic elasticity-mediated interaction plays a crucial role [92, 93, 94]. Particularly, when only a small fraction of particles undergo rearrangements, the system remarkably loses its ability to further accumulate elastic stress: As suggested in Ref. [95], the failure of amorphous solids occurs when only about $10\%$ of particles are involved in STZs. As for LER, its response is very similar to that of amorphous solids, as shown in figure 10 and Ref. [83]. Especially, the end of its elastic deformation is accompanied by the emergence of cage-jump spots that manifest STZs [83]. With these considerations, the observation shown in figure 12 (c) is understandable: During the elastic deformation stage of LER, cage jumps are relatively rare, and $R_{\mathrm{cj}}(t)$ grows slowly with $t$. As $R_{\mathrm{cj}}(t)$ exceeds about $10\%$, the elasticity cannot be well sustained any longer. Then, plastic yielding and disintegration successively occur, accompanied by overwhelming cage jumps that result in the rapid growth of $R_{\mathrm{cj}}(t)$.
@@ III.2 Steady shear
Under steady shear, both samples exhibit shear thinning, as shown in figure 1. It can be described by a phenomenological expression $\eta(\dot{\gamma})\approx\eta_{\mathrm{eq}}\big{/}\left[1+(\tau_{\eta}\dot{\gamma})^{\lambda}\right]$, where $\eta_{\mathrm{eq}}$ is the zero-shear viscosity, and $\tau_{\eta}$ is a parameter with the dimension of time [2, 45, 96]. As $\dot{\gamma}$ increases from null, $\eta$ gradually deviates from $\eta_{\mathrm{eq}}$ and finally behaves with a power law $\eta(\dot{\gamma})\sim\dot{\gamma}^{-\lambda}$. To understand this complicated nonlinear behavior, we employ the solid-liquid duality by expressing the shear stress $\sigma$ as:
@@
where the first term in the right-hand side of equation 5 represents the contribution from slow particles, and the second term represents the contribution from fast particles. We discuss the former first. $p_{\mathrm{slow}}$ is the fraction of slow particles. As demonstrated in section II.3, most slow particles are involved in the elastic stage of LERs. Thus, this contribution is given by the average elastic stress of LER. Here, $G_{\mathrm{eff}}$ is the effective shear modulus of LER, and $\tau_{\mathrm{el}}$ is the average time of duration of the elastic deformation of LER. In flow, an LER undergoes elastic loading, and then a sharp disintegration to unload. The factor $1/2$ accounts for the averaging effect of this saw-tooth-like loading-unloading process.
$\tau_{\mathrm{el}}$ and $G_{\mathrm{eff}}$ are needed to quantify the contribution from slow particles. $\tau_{\mathrm{el}}$ is evaluated as follows. As we discuss in the preceding subsection, the end of the elastic deformation stage of an LER is accompanied by the cage jumps of about $10\%$ of its constituent particles. Therefore, we determine $\tau_{\mathrm{el}}$ as the waiting time that $10\%$ of slow particles have relaxed according to the distribution of $\tau_{\mathrm{sh}}$ value.
We now turn our attention to the effective modulus of LER $G_{\mathrm{eff}}$. By “effective”, we mean that it encapsulates both elastic and plastic effects [71, 97]. To be specific, even in the elastic stage of LER, small amount of nonaffine displacements of constituent particles take place, as shown in figure 12 (c) and Ref. [83]. This plasticity slightly weakens the modulus of LER. Therefore, the modulus during the elastic stage can be written as $G_{\mathrm{eff}}=wG_{\mathrm{ini}}$, where $G_{\mathrm{ini}}$ is the initial shear modulus of LER, and $w$ ($w<1$) is a weakening factor reflecting this plasticity. To estimate $w$, we adopt a strategy that ties the variation in shear modulus to the neighbor loss process [71, 97]. The detail for calculating $w$ and $G_{\mathrm{eff}}$ is given in Appendix B. For all studied conditions, we find that $G_{\mathrm{eff}}\approx 0.8G_{\mathrm{ini}}$.
For fast particles, as we show in figure 10, they do not exhibit elastic deformation. Instead, such regions behave like viscous liquids. Considering that the average persistence time is found to be proportional to viscosity [62], we can estimate the viscosity of fast particles $\eta_{\mathrm{fast}}$ by the relation:
@@
where $\bar{\tau}_{\mathrm{LC,eq}}$ is the average of $\tau_{\mathrm{LC}}$ at equilibrium, and $\bar{\tau}_{\mathrm{sh,fast}}$ is the average of $\tau_{\mathrm{sh}}$ of fast particles at the given shear rate. Here, $\bar{\tau}_{\mathrm{sh}}$ can be approximated by $\bar{\tau}_{\mathrm{LC}}$ for fast particles. With equation 6, we obtain the second term in the right-hand side of equation 5.
Figure 13: In both (a) (representing sample A) and (b) (representing sample B), the viscosity estimated by equation 5, the viscosity directly measured from MD data, and the proportion of slow particles $p_{\mathrm{slow}}$ are given.
Figure 13 shows the results of viscosity calculated by equation 5 for two samples. The accurate results, shown in figure 1, are replotted here for comparison. The agreement between two results is remarkable. We also show the results of $p_{\mathrm{slow}}$ in figure 13. It is seen that $p_{\mathrm{slow}}$ is negligible at the Newtonian regime. As $\dot{\gamma}$ increases, $p_{\mathrm{slow}}$ becomes finite, meanwhile, noticeable shear thinning appears. As $\dot{\gamma}$ further increases, $p_{\mathrm{slow}}$ grows rapidly, and then saturates in the shear thinning regime. These observations clearly reveal the relation between DH and the onset of shear thinning: Due to the heterogeneous dynamics, different regions in a supercooled liquid respond to shear in different ways. Slow particles tend to aggregate into LERs. The accumulation of elastic stress and strain of LER, which is limited by thermal activation and shear-driven effects, cannot catch up with the growth of $\dot{\gamma}$. Thus, the viscosity contributed by slow particles shear-thins. As $\dot{\gamma}$ becomes larger, more particles are involved as slow particles, results in progressively prominent shear thinning. When $\dot{\gamma}$ is very large, most particles are involved in LERs. Consequently, the nonlinear rheology of the system, described by $\eta\propto\dot{\gamma}^{-\lambda}$, is dominated by the behavior of LER. As we demonstrate in Ref. [44, 83], in the shear-thinning regime, the characteristic length $\xi_{\mathrm{LER}}$ and strain $\gamma_{\mathrm{LER}}$ of LER, respectively, depend on $\dot{\gamma}$ by $\xi_{\mathrm{LER}}\propto\dot{\gamma}^{-v}$ and $\gamma_{\mathrm{LER}}\propto\dot{\gamma}^{\epsilon}$. Then, the exponent $\lambda$ is determined by LER through the scaling relation $\lambda=1-\epsilon=(D+1)v$, where $D$ is the dimension of the system.